bill533:

A line parallel to a triangle’s side splits one side into lengths of 9 and 3. The other side is split into lengths of 12 and x. What is the value of x? (Enter the number only)

bill533:

@563blackghost

jhonyy9:

 use the proportionality of corresponding sides

jhonyy9:

no proportionality of corresponding sides .... so how you ve got this result ?

bill533:

EF =

jhonyy9:

no you need calcule the value of x

bill533:

is it 4

jhonyy9:

9/3 = 12/x yes exactly

jhonyy9:

questions ?

bill533:

Did i had to divide by 9 9x/9 = 26/9 x = 4

jhonyy9:

no 26 you wan there 36

jhonyy9:

bc 36/9 = 4

jhonyy9:

ok. ?

bill533:

ok i got you but could you help me with more of my math questions

jhonyy9:

ok. ?

bill533:

ok

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 9/3 = 12/x cross multiply 9x = 3*12 9x = 36 divide both sides by 9 x = 36 / 9 x = 4 \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

ok - was my pleasure bye bye

bill533:

could you help me with more

jhonyy9:

post it on the open question column please

bill533:

yes sir